If dy/dx= x-x²/2x⁴ and y=2 when x=1 to express y in terms of x.

1
Expert's answer
2022-03-20T09:43:32-0400
"\begin{gathered}
\int \frac{x-x^{2}}{2 x^{4}} \mathrm{~d} x \\
=\int\left(\frac{1}{2 x^{3}}-\frac{1}{2 x^{2}}\right) \mathrm{d} x \\
=\frac{1}{2} \int \frac{1}{x^{3}} \mathrm{~d} x-\frac{1}{2} \int \frac{1}{x^{2}} \mathrm{~d} x
\end{gathered}
\\[4mm]
\text { Now solving: }\\[2mm]
\int \frac{1}{x^{3}} d x\\[2mm]

\text{Apply power rule:}\\[2mm]
\begin{aligned}
\int x^{\mathrm{n}} \mathrm{d} x=& \frac{x^{\mathrm{n}+1}}{\mathrm{n}+1} \text { with } \mathrm{n}=-3: \\
&=-\frac{1}{2 x^{2}}
\end{aligned}
\\[2mm]
\text{ Now solving:}\\[2mm]
\int \frac{1}{x^{2}} \mathrm{~d} x\\[2mm]
\text{ Apply power rule with } n =-2 :\\[2mm]
=-\frac{1}{x}\\[3mm]
\text{ Plug in solved integrals:}\\[3mm]
\begin{gathered}
\frac{1}{2} \int \frac{1}{x^{3}} \mathrm{~d} x-\frac{1}{2} \int \frac{1}{x^{2}} \mathrm{~d} x \\
=\frac{1}{2 x}-\frac{1}{4 x^{2}}
\end{gathered}
\\[2mm]
\text{ The problem is solved:}\\[2mm]
\begin{aligned}
& \int \frac{x-x^{2}}{2 x^{4}} \mathrm{~d} x \\
=& \frac{1}{2 x}-\frac{1}{4 x^{2}}+C
\end{aligned}
\\[3mm]
\text{ Simplifying the above, we have:}\\[2mm]
y=\frac{2 x-1}{4 x^{2}}+C"

At y=2 when x=1, we have:

"2=\frac{2 (1)-1}{4 (1)^{2}}+C\\[1.5mm]
2 = \frac{2 -1}{4 }+C\\[1.5mm]
2 = \frac{1}{4 }+C\\[1.5mm]
2 -\frac{1}{4 } = C\\[2mm]
\implies C = 1.75\\[4mm] 
\therefore y=\frac{2x-1}{4x^{2}}+1.75"